OpenStudy (anonymous):

Simplify 8 x3 / 2 x-3

OpenStudy (anonymous):

is this \[\frac{8\times 3}{2\times (-3)}\]?

OpenStudy (anonymous):

if so, you have one minus sign so answer is negative. 3 goes in to 3 once, 2 goes on to 8 4 times so answer is -4

OpenStudy (anonymous):

Thanks